OpenStudy (anonymous):

Factor completely p4 – 81

OpenStudy (anonymous):

I dont even get it, what times 81 twice then adds to p^4?

OpenStudy (cwrw238):

write it as (p^2)^2 - (9)^2

OpenStudy (cwrw238):

now apply rule for difference of 2 squares a^2 - b^2 = (a - b)(a + b)

OpenStudy (cwrw238):

do you follow that?

OpenStudy (anonymous):

(p – 3)^2(p + 3)^2

OpenStudy (anonymous):

yeah, thats my answer than right

OpenStudy (cwrw238):

a = p^2 and b = 9

OpenStudy (cwrw238):

(p^2)^2 - (9)^2 = (p^2 - 9)(p^2 + 9) - now you apply rule to p^2 - 9

OpenStudy (anonymous):

You can also brute force this if you like: